OpenStudy (anonymous):

What is -32-3/5f=-17?

OpenStudy (anonymous):

\[-32-\frac{3}{5}x=-17\]?

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

add 32 to both sides first

OpenStudy (anonymous):

Ok now I have 3/5x=15

OpenStudy (anonymous):

looks like you dropped the minus sign, a common but fatal mistake

OpenStudy (anonymous):

should be \[-\frac{3}{5}x=15\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now what

OpenStudy (anonymous):

multiply both sides by the reciprocal of \(-\frac{3}{5}\) which is \(-\frac{5}{3}\) to get \(x\) by itself

OpenStudy (anonymous):

i.e. write \[x=15\times \left(-\frac{3}{5}\right)\] cancel first , multiply last

OpenStudy (anonymous):

So it's -45?

OpenStudy (anonymous):

No it is -9.. when u multiply 15 by -3, u get -45.. divide -45 by 5 and u get -9